# From Counterfactuals to Trees:Competitive Analysis of Model Extraction Attacks

### v1

@@ 3.2 Tree Reconstruction Attack algorithm
The proofs of Propositions 3.6 and 3.7 are provided in Appendix A. Proposition 3.6 provides the competitive ratio achieved by the TRA algorithm, while Proposition 3.7 establishes that the choice of query position does not affect the competitive ratio for any divide-and-conquer algorithm iteratively partitioning the input space. Together, these propositions demonstrate that TRA not only offers a competitive approach to tree reconstruction under various query selection strategies but also sets a theoretical benchmark that cannot be surpassed by other divide-and-conquer-based methods. This underscores the effectiveness of TRA in efficiently reconstructing axis-parallel decision boundary models.
Anytime Behavior.
Since the query budget may not be known in advance, and may vary depending on the application, the ability of an extraction attack to operate in an anytime manner is crucial. It means that it can be stopped at any iteration while still producing a decision tree classifier—though it may not yet be fully functionally equivalent. Specifically, whenever we split the input space (line 9 of Algorithm 1), we assign provisional labels to the resulting subspaces: one inherits the query’s label, and the other adopts the counterfactual’s label. This design ensures the algorithm can output a classifier at any stage of execution. Note that the exploration strategy of Algorithm 1 depends on the ordering of the priority queue $\mathcal{Q}$. In practice, preliminary experiments suggested good anytime performance using breadth-first search (BFS), which distributes the exploration evenly. While the design of alternative best-first search strategies is a promising direction to enhance TRA’s anytime performance, it is worth noting that the exploration strategy does not affect the number of queries needed by TRA to fully reconstruct the target model. Indeed, it does not modify the iterative partitioning of the input space performed by the divide-and-conquer algorithm.
@@ 4.1 Experimental Setup
Datasets.
We use six binary classification datasets, selected from related works on model extraction attacks (Aïvodji et al., 2020; Wang et al., 2022; Tramèr et al., 2016) and encompassing a variety of feature types, dimensionalities, and classification tasks, as summarized in Table 1. More precisely, we consider the COMPAS dataset (Angwin et al., 2016), as well as the Adult Income (Adult), Default of Credit Card Clients (Credit Card), German Credit and Student Performance (SPerformance) datasets from the UCI repository (Dua & Graff, 2017). Categorical features are one-hot encoded, while numerical, discrete (ordinal) and binary ones are natively handled by both tree building procedures and reconstruction attacks. Each dataset is partitioned into training, validation, and test sets with proportions of 60%, 20%, and 20%, respectively.
@@
We train two types of tree-based target models implemented in the scikit-learn library (Pedregosa et al., 2011): decision trees and random forests. For decision trees, we experiment with varying max_depth parameters ranging from 4 to 10, as well as unbounded trees (max_depth set to None). The experiments on random forests focus on the COMPAS dataset, employing different numbers of estimators to assess scalability and robustness. To prevent overfitting, we utilize the validation set for hyperparameter tuning and apply cost-complexity pruning where applicable. Detailed training procedures and hyperparameters configurations are elaborated in Appendix B.1.
Baselines.
@@
We benchmark TRA against three state-of-the-art model extraction attacks. First, PathFinding is the only functionally equivalent model extraction attack against decision trees. While it does not rely on counterfactual examples, it assumes access to a leaf identifier indicating in which leaf of the target decision tree the query example falls (Tramèr et al., 2016). It is thus not applicable to random forests. Second, CF (Aïvodji et al., 2020) leverages counterfactual explanations to train a surrogate model mimicking the target one. Third, DualCF (Wang et al., 2022) enhances the CF approach by additionally computing the counterfactuals of the counterfactuals themselves, which improves fidelity (Wang et al., 2022). For each baseline, we adapt the number of queries relative to the complexity of the target model. Specifically, CF and DualCF are allocated a query budget of $50$ times the number of nodes in the target decision tree, while PathFinding is configured with an $\epsilon=10^{-5}$ to achieve approximate functional equivalence.
@@
We evaluate three surrogate model variants for the two surrogate-based baselines CF and DualCF: a multilayer perceptron (MLP), and two models from the same hypothesis class as the target model (i.e., a decision tree or a random forest) – one of them sharing the exact same hyperparameters, and the other one using default hyperparameter values. These three surrogates correspond to three different knowledge for the adversary: knowledge of the hypothesis class of the target model, of its hyperparameters, or none of them. Both CF and DualCF were originally tested against heuristic-based counterfactual explanations using the DiCE (Mothilal et al., 2020) algorithm. To assess the impact of explanation optimality on the attack’s performance, and to ensure fair comparisons, we run CF and DualCF using both DiCE and the OCEAN framework (Parmentier & Vidal, 2021), the later providing optimal counterfactual explanations. In a nutshell, OCEAN formulates the search of counterfactual examples as an optimization problem, modeled and solved using a mixed-integer linear programming solver. Our implementation of TRA relies on OCEAN as the counterfactual oracle $\mathcal{O}_{d}$.
Evaluation.
We employ two metrics to quantify the effectiveness of each model extraction attack. First, fidelity quantifies the success of the performed extraction attack, by measuring the proportion of inputs on a given set for which the extracted model’s predictions match those of the target model. Second, the number of queries indicates the attack’s efficiency, as it counts the total number of queries made to the target model’s oracle during the extraction process.
@@
All experiments are run on a computing platform, over a set of homogeneous nodes using Intel Platinum 8260 Cascade Lake @ 2.4GHz CPU. Each run is allocated 4 threads using up to $4$GB of RAM each (multi-threading is mainly used by the OCEAN optimal counterfactual oracle). We repeat each experiment five times, using five different random seeds, and report average values. The source code to reproduce all our experiments and figures is accessible at https://github.com/vidalt/Tree-Extractor, under a MIT license.
@@ 4.2 Results
Figure 3: Anytime performance of all the considered model extraction attacks against decision trees. We report results for two datasets and retain the best configuration for the surrogate-based attacks CF and DualCF.
Figure 4: Anytime performance of the considered model extraction attacks against random forests, on the COMPAS datasets.
@@
Figure 5: Performance of the PathFinding and TRA functionally equivalent model extraction attacks against decision trees. We report results for two datasets where each point represent the number of queries needed to fully reconstruct the trees.
Fidelity measurements.
We consider two fidelity values, measured either on a dataset of $3,000$ examples uniformly sampled over the input space or on a test set. Since both measurements provide the same trends, we hereafter focus on the later, which better characterizes the ability of the extracted models to mimic the target ones over the whole input space. For the sake of completeness, we also report fidelity results measured on a test set for our experiments using random forest target models in Appendix B.3.
Tuning surrogate-based model extraction attacks.
In Appendix B.2, we provide the detailed results of our model extraction experiments using the CF and DualCF attacks. We perform a comprehensive comparison of the impact of the type of surrogate and counterfactual oracle used. Our results demonstrate that fitting surrogate models of the same hypothesis class facilitates the extraction of both decision trees and random forests. However, knowledge of their hyperparameters does not provide any advantage to either attack. Finally, the non-optimal counterfactuals provided by DiCE lead to better fidelity results than the optimal ones computed by OCEAN for these two attacks. This can be explained by their heuristic nature, which leads to the building of more diverse counterfactuals, not necessarily lying next to a decision boundary. For the remaining of this section, we retain only the best performing configuration for both CF and DualCF, which consists in training a surrogate model of the same hypothesis class using DiCE counterfactuals.
TRA outperforms existing approaches in terms of anytime fidelity to extract tree-based models.
Figure 3 presents the average fidelity of surrogate models obtained from the four studied model extraction attacks on decision trees, as a function of the number of queries, for the Adult and COMPAS datasets. Results for additional datasets, provided in Figure 11 (Appendix B.4), exhibit the same trends. Across all cases, TRA consistently achieves higher fidelity for any fixed query budget and converges orders of magnitude faster to perfect fidelity, while also certifying functional equivalence—unlike CF and DualCF. The anytime performance of the considered model extraction attacks against random forests (excluding PathFinding, which is not applicable to ensemble models) is shown in Figure 4 for the COMPAS dataset. Again, TRA outperforms existing methods, producing more accurate surrogate models with fewer queries and converging to perfect fidelity significantly faster.
TRA exhibits state-of-the-art performances for functionally equivalent extraction of decision trees.
Figure 5 shows the number of queries required by PathFinding and TRA to achieve functionally equivalent model extraction, plotted against the number of nodes in the target models. Results for the Adult and COMPAS datasets are presented, with additional datasets provided in Figure 12 (Appendix B.4). TRA consistently requires orders of magnitude fewer queries than PathFinding to reconstruct the target models with perfect fidelity.
TRA theoretically and empirically outperforms existing approaches to extract random forests.
@@
Figure 4 presents the average fidelity of the three considered model extraction attacks against random forests, plotted against the number of performed queries for the COMPAS dataset. The results show that TRA achieves higher fidelity with fewer queries to the counterfactual oracle and converges significantly faster to perfect fidelity. Moreover, TRA is the only attack that certifies functional equivalence for tree ensemble target models. Additional results in Figure 13 (Appendix B.4) indicate that TRA scales efficiently with the size of the target random forest, as the number of required queries grows sub-linearly with the total number of nodes. This behavior arises because large forests with many trees may introduce redundancies, allowing the extracted model to be represented with perfect fidelity as a more compact decision tree (Vidal & Schiffer, 2020).
@@ 5 Related Works
The flourishing literature on privacy in machine learning encompasses a wide variety of inference attacks, each considering different setups and objectives (Cristofaro, 2020; Rigaki & García, 2024). This paper focuses on model extraction attacks (Tramèr et al., 2016), which aim at reconstructing the decision boundary of a black-box target model as accurately as possible, given a prediction API. As highlighted in recent surveys (Gong et al., 2020; Oliynyk et al., 2023), numerous attacks have been proposed in recent years, targeting a broad spectrum of hypothesis classes. Hereafter, we focus on those specifically targeting tree-based models or exploiting counterfactual explanations.
Tramèr et al. (2016) propose a functionally equivalent model extraction attack specifically targeting regression or decision trees: PathFinding. Their approach assumes that each query reply contains a unique identifier for the associated leaf. In a nutshell, PathFinding aims at identifying the decision boundaries of each leaf in the target tree by varying the values of each feature. While effective, this method requires a large number of queries, though partial input queries can sometimes mitigate this overhead.
While many recent have focused on generating counterfactual explanations (Guidotti, 2024), they have also been shown to facilitate privacy attacks (Pawelczyk et al., 2023). Aïvodji et al. (2020) introduce CF, a model extraction attack that leverages counterfactual explanations. Their approach constructs a labeled dataset by querying both predictions and counterfactuals from the target model, which is then used to train a surrogate. Wang et al. (2022) extend this method with DualCF, which improves fidelity by additionally querying the counterfactuals of the counterfactual explanations. Dissanayake & Dutta (2024) employ polytope theory to show that a sufficient number of optimal counterfactual explanations can approximate convex decision boundaries. They propose a model extraction attack against locally Lipschitz continuous models, with fidelity guarantees dependent on the Lipschitz constants of the target and surrogate models. However, functional equivalence cannot be strictly certified, and as the authors acknowledge, these guarantees do not apply to decision tree-based models, which lack local Lipschitz continuity and convexity. Finally, while beyond the scope of this paper, other explanation-based model extraction attacks have been explored, including those relying on gradient-based (Milli et al., 2019; Miura et al., 2024) and other feature-based methods (Öksüz et al., 2024).
@@ 6 Conclusions and Discussion
We introduced the first functionally equivalent model extraction attack against decision trees and tree ensembles, leveraging optimal counterfactual explanations. In addition to its rigorous functional equivalence guarantee, the proposed method achieves higher fidelity than prior approaches while requiring fewer queries. We also leveraged well-established tools from online discovery to enable a formal analysis of model extraction, drawing an analogy between the two fields. We illustrated the applicability and relevance of this analysis by providing bounds on our attack’s efficiency compared to the best achievable strategy, relying on the notion of competitive ratio. This perspective is essential for formally characterizing and comparing model extraction attacks.
Our study demonstrates that optimal counterfactual explanations can be systematically exploited to reconstruct tree ensembles via query APIs, as they inherently reveal decision boundaries. This raises significant concerns, especially as explainability is increasingly mandated by regulations. In many real-world applications, counterfactual explanations serve as a natural mechanism to meet transparency requirements by providing recourse information.
The research perspectives connected to our work are numerous. First, we believe that competitive analysis provides a valuable foundation for studying model extraction attacks, and future work should adopt the same lenses to evaluate other target models. Besides this, both the algorithms and their theoretical bounds could be refined. Improving input space exploration while mitigating worst-case query complexity is a key direction, including strategies such as dynamically reordering TRA’s priority queue to avoid worst-case scenarios, or adopting completely different exploration methods. Finally, investigating the impacts of privacy-preserving mechanisms for counterfactual explanations (Vo et al., 2023) on attack’s success is a crucial direction towards conciliating trustworthiness and privacy through ML explainability APIs.
Impact Statement
As a result, MLaaS platforms increasingly integrate such explainability tools into their APIs. While these explanations enhance user trust, they also expose a new attack surface to malicious entities by revealing additional model information. In this work, we theoretically and empirically demonstrate that optimal counterfactual explanations of decision trees and tree ensembles can be exploited to conduct efficient model extraction attacks. These results highlight the critical tension between transparency and the protection of model integrity and intellectual property. By identifying and quantifying these vulnerabilities, we highlight the risks of releasing model explanations without a thorough security assessment. Our research establishes a benchmark for evaluating method safety, advocating for the development of privacy-preserving approaches to explainability.

### v2

@@ 3.2 Tree Reconstruction Attack algorithm
The proofs of Propositions 3.6 and 3.7 are provided in Appendix A. Proposition 3.6 provides the competitive ratio achieved by the TRA algorithm, while Proposition 3.7 establishes that the choice of query position does not affect the competitive ratio for any divide-and-conquer algorithm iteratively partitioning the input space. These propositions demonstrate that TRA not only offers a competitive approach to tree reconstruction under various query selection strategies, but also sets a theoretical benchmark that cannot be surpassed by other divide-and-conquer-based methods.
Anytime Behavior. Since the query budget is often not known in advance and may vary depending on the target model, it is crucial for an extraction attack to operate in an anytime fashion—that is, to produce a usable classifier even if interrupted before completion. TRA satisfies this property by assigning provisional labels at each input space split (line 9 of Algorithm 1): one subregion inherits the query’s label, the other the counterfactual’s. This guarantees that a valid decision tree classifier is available at any point during execution. The quality of intermediate classifiers depends on the ordering of the priority queue $\mathcal{Q}$. In practice, preliminary experiments suggested good anytime performance using breadth-first search (BFS), which distributes the exploration evenly. While the design of alternative priority strategies is a promising research direction to enhance TRA’s anytime performance, the choice of exploration order does not impact the total number of queries required for exact reconstruction —this is solely determined by the algorithm’s divide-and-conquer structure.
@@ 4.1 Experimental Setup
Datasets. We use five binary classification datasets, selected from related works on model extraction attacks [Aïvodji et al., 2020, Wang et al., 2022, Tramèr et al., 2016] and encompassing a variety of feature types, dimensionalities, and classification tasks, as summarized in Table 1. More precisely, we consider the COMPAS dataset [Angwin et al., 2016], as well as the Adult Income (Adult), Default of Credit Card Clients (Credit Card), German Credit and Student Performance (SPerformance) datasets from the UCI repository [Dua and Graff, 2017]. Categorical features are one-hot encoded, while numerical, discrete (ordinal) and binary ones are natively handled by both tree building procedures and reconstruction attacks. Each dataset is partitioned into training, validation, and test sets with proportions of 60%, 20%, and 20%, respectively.
@@
Training the target tree-based models. We train two types of tree-based target models implemented in the scikit-learn library [Pedregosa et al., 2011]: decision trees and random forests. For decision trees, we experiment with varying max_depth parameters ranging from 4 to 10, as well as trees without maximum depth constraint (max_depth set to None). The random forests experiments focus on the COMPAS dataset, employing different numbers of trees (5, 25, 50, 75 and 100) to assess scalability and robustness. To prevent overfitting, we utilize the validation set for hyperparameter tuning and apply cost-complexity pruning where applicable. All the details of training procedures and hyperparameters configurations are discussed in Appendix B.1.
@@
Baselines. We benchmark TRA against three state-of-the-art model extraction attacks. First, PathFinding [Tramèr et al., 2016] is the only functionally equivalent model extraction attack against decision trees. While it does not rely on counterfactual examples, it assumes access to a leaf identifier indicating in which leaf of the target decision tree the query example falls. It is thus not applicable to random forests. Second, CF [Aïvodji et al., 2020] leverages counterfactual explanations to build a labeled attack set and train a surrogate model mimicking the target one. Third, DualCF [Wang et al., 2022] enhances the CF approach by additionally computing the counterfactuals of the counterfactuals themselves, which has been shown to improve fidelity. We adapt the number of queries (which must be pre-fixed for both CF and DualCF) to the complexity of the target model as it is set to $50$ times the number of nodes in the target decision tree. PathFinding is configured with $\epsilon=10^{-5}$ (pre-fixed precision of the retrieved splits) to achieve approximate functional equivalence.
@@
We evaluate three surrogate model variants for CF and DualCF: a multilayer perceptron (MLP), and two models from the same hypothesis class as the target model (i.e., a decision tree or a random forest), one of them sharing the exact same hyperparameters, and the other using default hyperparameter values. These variants reflect different levels of adversarial knowledge: the hypothesis class, the exact hyperparameters, or neither. Both CF and DualCF were originally evaluated using heuristic counterfactual explanations from the DiCE [Mothilal et al., 2020] algorithm. To assess the impact of explanation optimality on the attack’s performance, and to ensure fair comparisons, we tested these baselines using either DiCE or the OCEAN framework [Parmentier and Vidal, 2021], which formulates counterfactual search as a mixed-integer linear program and guarantees optimality.
Complete experimental results across all configurations of CF and DualCF are reported in Appendix B.2. They demonstrate that fitting surrogate models of the same hypothesis class facilitates the extraction of both decision trees and random forests. However, knowledge of their hyperparameters does not provide any advantage to either attack. Finally, the non-optimal counterfactuals provided by DiCE lead to better fidelity results than the optimal ones computed by OCEAN for these two attacks. This can be explained by their heuristic nature, which leads to the building of more diverse counterfactuals, not necessarily lying next to a decision boundary. In the next section, we display results only for the best-performing configuration for both CF and DualCF, achieved by training a surrogate model of the same hypothesis class and using DiCE counterfactuals.
Evaluation. We assess each model extraction attack using two metrics: fidelity and number of queries made to the prediction and counterfactual oracle API during the attack. Fidelity measures the proportion of inputs for which the extracted model agrees with the target model, quantifying attack success. We compute fidelity over 3000 points sampled uniformly from the input space, providing a broad evaluation across the entire feature domain. For completeness, we also report fidelity measured on a test set (i.e., drawn from the data distribution) for our experiments using random forest target models, in Appendix B.3, which shows the same performance trends.
Counterfactual oracle. As discussed in Section 3.2, TRA requires the use of a locally optimal counterfactual oracle. Since global optimality is a sufficient condition, we use the popular OCEAN oracle for simplicity in our main experiments with TRA. In Appendix C, we report the performance of TRA using a simple heuristic oracle that produces locally optimal explanations. These experiments show that the choice of oracle (optimal or heuristic) has minimal impact on reconstruction performance.
@@
All experiments are run on a computing cluster with homogeneous nodes using Intel Platinum 8260 Cascade Lake @ 2.4GHz CPU. Each run uses four threads and up to $4$GB of RAM each (multi-threading is only used by the OCEAN oracle). We repeat each experiment five times with different random seeds and report average values. The source code to reproduce all our experiments and figures is accessible at https://github.com/vidalt/Tree-Extractor, under a MIT license.
@@ 4.2 Results
Result 1. TRA outperforms existing approaches in terms of number of queries and anytime fidelity to extract decision trees. Figure 3 presents the average fidelity of surrogate models obtained from the four studied model extraction attacks on decision trees, as a function of the number of queries, for the Adult and COMPAS datasets. Here, each point on a curve is the mean fidelity over all 40 extraction tasks (eight tree depths $\times$ five seeds), so when it reaches 1.00, all target trees were perfectly reconstructed. Results for additional datasets, provided in Figure 11 (Appendix B.4), exhibit the same trends. Across all cases, TRA consistently achieves higher fidelity for any fixed query budget and converges to perfect fidelity orders of magnitude faster. Notably, unlike CF and DualCF, TRA and PathFinding provide formal guarantees of functional equivalence.
Result 2. TRA achieves state-of-the-art performance for functionally equivalent extraction of decision trees. Figure 4 shows the number of queries required by PathFinding and TRA to achieve functionally equivalent model extraction, as a function of the number of nodes in the target models. Results for the Adult and COMPAS datasets are presented, with additional datasets provided in Figure 12 (Appendix B.4). TRA consistently requires orders of magnitude fewer queries than PathFinding to reconstruct the target models with perfect fidelity.
@@
Figure 3: Anytime performance of all the considered model extraction attacks against decision trees.
@@
Figure 4: Performance of the functionally equivalent model extraction attacks against decision trees. We report the number of queries required to fully reconstruct the trees as a function of their size.
Figure 5: Anytime performance of the considered model extraction attacks against random forests, on the COMPAS dataset.
@@
Result 3. TRA theoretically and empirically outperforms existing approaches to extract random forests. Figure 5 presents the average fidelity of the three considered model extraction attacks against random forests, plotted against the number of performed queries for the COMPAS dataset. The results show that TRA achieves higher fidelity with fewer counterfactual queries and converges significantly faster to perfect fidelity. Moreover, TRA is the only attack that certifies functional equivalence for tree ensembles. Additional results in Figure 13 (Appendix B.4) indicate that TRA scales efficiently with the size of the target random forest, as the number of required queries grows sub-linearly with the total number of nodes. This is due to structural redundancies in large forests, allowing the extracted model to be represented with perfect fidelity as a more compact decision tree [Vidal and Schiffer, 2020].
Result 4. TRA is effective with either globally or locally optimal counterfactuals. We report in Appendix C results from experiments where TRA is run with a simple oracle generating heuristic counterfactuals that are only locally optimal. The results show that TRA’s performance is largely unaffected by the lack of global optimality. In some cases, locally optimal counterfactuals even improve early-stage (anytime) performance by introducing greater diversity in the explored input space. These results highlight the practical applicability of TRA in real-world scenarios, where an attacker—given access to any counterfactual oracle—can use simple query access to the target model to iteratively refine counterfactuals via line search, making them locally optimal and suitable to conduct an efficient model extraction attack.
@@ 5 Related Works
The flourishing literature on privacy in machine learning encompasses a wide variety of inference attacks, considering different setups and objectives [Cristofaro, 2020, Rigaki and García, 2024]. This paper focuses on model extraction attacks [Tramèr et al., 2016], which aim at reconstructing the decision boundary of a black-box target model as accurately as possible, given a prediction API. As highlighted in recent surveys [Gong et al., 2020, Oliynyk et al., 2023], various attacks have been proposed in recent years, targeting a broad spectrum of hypothesis classes. Hereafter, we focus on those targeting axis-parallel decision boundary models or exploiting counterfactual explanations.
Tramèr et al. [2016] propose the only functionally equivalent model extraction attack targeting regression or decision trees: PathFinding. It assumes that each query reply contains a unique identifier for the associated leaf. In a nutshell, PathFinding identifies the decision boundaries of each leaf in the target tree by varying the values of each feature. While effective, this method requires a large number of queries, though partial input queries can sometimes mitigate this overhead.
While many recent works have focused on generating counterfactual explanations [Guidotti, 2024], these explanation techniques have also been shown to facilitate privacy attacks [Pawelczyk et al., 2023]. Aïvodji et al. [2020] introduce CF, a model extraction attack that leverages counterfactual explanations. Their approach constructs a labeled dataset by querying both predictions and counterfactuals from the target model, which is then used to train a surrogate. Wang et al. [2022] extend this method with DualCF, which improves fidelity by additionally querying the counterfactuals of the counterfactual explanations. Dissanayake and Dutta [2024] employ polytope theory to show that a sufficient number of optimal counterfactual explanations can approximate convex decision boundaries. They propose a model extraction attack against locally Lipschitz continuous models, with fidelity guarantees dependent on the Lipschitz constants of the target and surrogate models. However, functional equivalence cannot be strictly certified, and as the authors acknowledge, these guarantees do not apply to decision tree-based models, which lack local Lipschitz continuity and convexity.
Finally, while beyond the scope of this paper, other explanation-based model extraction attacks have been explored, including those relying on gradient-based [Milli et al., 2019, Miura et al., 2024] and other feature-based methods [Öksüz et al., 2024].
@@ 6 Conclusions and Discussion
We introduced the first functionally equivalent model extraction attack against decision trees and tree ensembles, leveraging locall optimal counterfactual explanations. In addition to its rigorous functional equivalence guarantee, the proposed method achieves higher fidelity than prior approaches while requiring fewer queries. We also leveraged well-established tools from online discovery to enable a formal analysis of model extraction, drawing an analogy between the two fields. We illustrated the applicability and relevance of this analysis by providing bounds on our attack’s efficiency compared to the best achievable strategy, relying on the notion of competitive ratio. This perspective is essential for formally characterizing and comparing model extraction attacks.
Our study demonstrates that optimal counterfactual explanations can be systematically exploited to reconstruct tree ensembles via query APIs, as they inherently reveal decision boundaries. This raises significant concerns, especially as explainability is increasingly mandated by regulations. In many real-world applications, counterfactual explanations serve as a natural mechanism to meet transparency requirements by providing recourse information. We discuss the broader societal impacts of our work in Appendix D.
The research perspectives connected to our work are numerous. First, we believe that competitive analysis provides a valuable foundation for studying model extraction attacks, and future work should adopt the same lenses to evaluate other target models. Besides this, both the algorithms and their theoretical bounds could be refined. Improving input space exploration while mitigating worst-case query complexity is a key direction, including strategies such as dynamically reordering TRA’s priority queue to avoid worst-case scenarios, or adopting completely different exploration methods (beyond divide-and-conquer-based algorithms). Finally, investigating the impacts of privacy-preserving mechanisms for counterfactual explanations [Vo et al., 2023] on model extraction attacks’ success is a crucial direction towards conciliating trustworthiness and privacy through ML explainability APIs.
Acknowledgements
This research was enabled by support provided by Calcul Québec and the Digital Research Alliance of Canada, as well as funding from the SCALE AI Chair in Data-Driven Supply Chains.